OpenStudy (anonymous):

Jennie purchased 5 of the 200 tickets at the raffle. There will be two drawings for a prize. The probability that Jennie will win both of the prizes is 1 out of ____.

OpenStudy (anonymous):

5/200 * 4/199, assuming they don't put the first drawn ticket back after drawing.

myininaya (myininaya):

\[\frac{20}{200}*\frac{1}{199}=\frac{1}{10}*\frac{1}{199}=\frac{1}{1990}\]